OpenStudy (anonymous):

The value of y varies directly with x, and y = 12 when x = 8. Find y when x = 15.

OpenStudy (texaschic101):

y = kx (Direct variation formula) Now sub 12 in for y and 8 in for x 12 = k(8) 12/8 = k 3/2 = k now , use the same formula, and sub 15 in for x y = 3/2(15) y = 45/2 or 22 1/2 or 22.5 or this way.... 8/12 = 15/x (set it up as a proportion) cross multiply.. (8)(x) = (12)(15) 8x = 180 x = 180/8 x = 22 1/2 or 22.5